OpenStudy (anonymous):

Brittany is making a model boat and needs 4 pieces of string to rig the sails. Each piece of string needs to be 4/5 foot. She decides to cut the 4 pieces from a longer piece of string that is 15.2 feet in length. How much string will remain after Brittany cuts the 4 pieces of string?Select one of the options below as your answer: A. 5 feet B. 12 feet C. 18.4 feet D. 10.4 feet

Directrix (directrix):

4/5 foot = .8 foot. The string is 15.2 feet in length and Brittany is cutting 4 pieces of length .8 from it. So, 15.2 - 4(.8) = 15.2 - 3.2 = 12 feet which is option B.